OpenStudy (anonymous):

A colony of bacteria is growing exponentially according to the function below, where t is in hours. How many hours will it take for the population to grow to 200. Round to nearest nteger B(t)=4*e^0.8t

OpenStudy (anonymous):

@dpaInc how did you get 8=e^(0.8t)

OpenStudy (anonymous):

divide 4 on both sides to get: 50=e^(0.8t) ln50=.08t (ln50)/(0.8)=t there... that's better...:)

OpenStudy (anonymous):

so im left with 4.89? @dpaInc

OpenStudy (anonymous):

yea.. that's what got too...